CURVATURE PROPERTIES

OF SOME HYPERSURFACES IN SPACES OF CONSTANT CURVATURE

Ryszard Deszcz and Małgorzata Głogowska

Abstract. We present curvature properties of pseudosymmetry type of some hypersurfaces in semi-Riemannian spaces of constant curvature. Results related to this talk are contained in the following papers: [1], [2], [3] and [4].

References

[1] R. Deszcz and M. Głogowska, Some examples of nonsemisymmetric Ricci-semisymmetric hypersurfaces, Colloq. Math., in print.

[2] R. Deszcz and M. Głogowska, On nonsemisymmetric Ricci-semisymmetric warped product hypersurfaces, Dept. Math., Agricultural Univ. Wrocław, Report No. 90, 2000.

[3] R. Deszcz, M. Głogowska, M. Hotloś, and L. Verstraelen, On some generalized Einstein metric conditions on hypersurfaces in semi-Riemannian space forms, Dept. Math., Agricultural Univ. Wrocław, Report No. 99, 2001.

[4] M. Głogowska, On a curvature characterization of Ricci-pseudosymmetric hypersurfaces, Dept. Math., Agricultural Univ. Wrocław, Report No. 97, 2001.

Ryszard Deszcz and Małgorzata Głogowska Department of Mathematics, Agricultural University of Wrocław ul. Grunwaldzka 53, PL - 50-357 Wrocław, POLAND E-mail: rysz@ozi.ar.wroc.pl and mglog@ozi.ar.wroc.pl